find the equation of the line with slope 5/7 and passing through (1 -6)
Answer by lwsshak3(11628)

find the equation of the line with slope 5/7 and passing through (1 -6)
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given line:
equation:
y=5x/7+b
solve for b using coordinates of given point on the line(1,-6)
-6=5*1/7+b
b=-42/7-5/7=-47/7
equation: y=5x/7-47/7